OpenStudy (anonymous):

I NEED SOME EMERGENCY HELP WILL MEDAL AND FAN PLEASE!!!! Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = #2ndicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4) #3ndicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4) #4Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (kropot72):

What is the value of m for the first part? You appear to have missed typing it.

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

sorry

OpenStudy (kropot72):

You need to substitute the given values into the equation y = mx + b and rearrange into standard form.

OpenStudy (kropot72):

This will let you find the value of b which is the y-intercept.

OpenStudy (anonymous):

okay so 5x-2y=2/3?

OpenStudy (anonymous):

wait i think im off

OpenStudy (anonymous):

what am i doing wrong?

OpenStudy (kropot72):

Substitute the given values into y = mx + b. Then solve to find the value of b. Can you do that as a first step?

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

this is what it looks like _x__y__

OpenStudy (kropot72):

y = -2, x = 5 and m = 2/5 Plug these values into y = mx + b What do you get?

OpenStudy (anonymous):

do i x 2/5 and 5?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

that would be 2

OpenStudy (kropot72):

So we should have -2 = 2 + b which needs to be solved to find the value of b.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ugh i dont know do we x 5 and 2?

OpenStudy (kropot72):

What do you get for the value of b?

OpenStudy (anonymous):

10? *I say very unsure*

OpenStudy (kropot72):

Can you solve -2 = 2 + b to find the value of b?

OpenStudy (anonymous):

oh wait i felt like it was harder then it was i hope would b be -4?

OpenStudy (anonymous):

so it would be -2=2+-4?

OpenStudy (kropot72):

Yes, b = -4. So we now have \[\large y=\frac{2}{5}x-4\] which needs to be rearranged into standard form ax + by = c

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

whats a?

OpenStudy (kropot72):

First interchange the left and right sides to give \[\large \frac{2}{5}x-4=y\] Then rearrange to give \[\large \frac{2}{5}x-y=4\] Now you need to remove the fraction by multiplying both sides by 5.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

hahaha hold i thats so wrong

OpenStudy (anonymous):

i hate math

OpenStudy (anonymous):

10x-y=4?

OpenStudy (kropot72):

\[\large 5(\frac{2}{5}x-y)=4\times5\] which becomes \[\large 2x-5y=20\]

OpenStudy (anonymous):

oh lol i feel stupid now

OpenStudy (anonymous):

thank you soooo much i think now i can figure out the rest now

OpenStudy (kropot72):

You're welcome :)